A man word six weeks on the job. His son worked one third as long. The sons weekly salary was 3/5 that of his father. If they earned a total of $132 what were their weekly salaries.
Answer by ewatrrr(24785)

(6wk)S + (2wk)(3/5)S = 132
Multiply thru by 5
30S + 6S = 660
S = 660/36 = $18.33, father's. Son $12.22